# Global Compression Commander: Plug-and-Play Inference Acceleration for High-Resolution Large Vision-Language Models

### v5

@@ Appendix A Theoretical Complexity Analysis
Figure 9: Detailed analysis of content-agnostic positional bias. LLM attention-guided methods (e.g., FastV (Chen et al. 2024b), PDrop (Xing et al. 2025)) assign disproportionately high scores (bars) to tokens in later positions, regardless of their actual semantic content or input order (second row: sequential crop input; third row: reverse crop input).
@@
Appendix D Detailed Experimental Settings
@@
  • MMStar (Chen et al. 2024a): Contains 1,500 challenging samples across 6 core capabilities (coarse perception, fine-grained perception, instance interaction, logical reasoning, science & technology, mathematics) designed to evaluate vision-indispensable reasoning without relying on shortcuts.
@@
  • FastV (Chen et al. 2024b) performs one-time pruning after the selected LLM layer based on attention weights between vision tokens and the last token.
@@
Appendix E Additional Experiments
@@
E.1 Comparisons on LLaVA-NeXT-13B
@@
E.2 Results on LLaVA-OneVision-0.5B
@@
E.3 Sensitivity Analysis of Hyper-parameters
@@
E.4 Detailed Efficiency Analysis
@@
E.5 More Visualizations of Token Compression
@@
Appendix F Algorithm Illustration

### v6

@@ 6 Conclusion
Acknowledgments
This work was supported in part by the Chengdu Science and Technology Program (No. 2025-YF12-00006-RC) , Police Integration Computing Key Laboratory of Sichuan Province (No. JWRH202502002), and the Open Fund of Key Laboratory of the Ministry of Education on Artificial Intelligence in Equipment (No. 2024-AAIE-KF04-03).
@@ Appendix A Theoretical Complexity Analysis
Figure 9: Detailed analysis of content-agnostic positional bias. LLM attention-guided methods (e.g., FastV (Chen et al. 2024a), PDrop (Xing et al. 2025)) assign disproportionately high scores (bars) to tokens in later positions, regardless of their actual semantic content or input order (second row: sequential crop input; third row: reverse crop input).
@@
Appendix D Extension to Video Understanding
Appendix E Detailed Experimental Settings
@@
  • MMStar (Chen et al. 2024b): Contains 1,500 challenging samples across 6 core capabilities (coarse perception, fine-grained perception, instance interaction, logical reasoning, science & technology, mathematics) designed to evaluate vision-indispensable reasoning without relying on shortcuts.
@@
  • FastV (Chen et al. 2024a) performs one-time pruning after the selected LLM layer based on attention weights between vision tokens and the last token.
@@
Appendix F Additional Experiments
@@
Method MVBench LongVideoBench MLVU VideoMME Average
Overall Short Medium Long
Upper Bound, 6272 Tokens
LLaVA-OneVision-7B 56.9 56.4 63.0 58.6 70.3 56.6 48.8 100.0%
Ratio=30%, Retain 1882 Tokens
DyCoke(CVPR25) 56.6 54.7 60.3 56.1 67.1 54.6 46.6 96.5%
Ratio=25%, Retain 1568 Tokens
FastV(ECCV24) 55.5 53.3 59.6 55.3 65.0 53.8 47.0 95.0%
PDrop(CVPR25) 55.3 51.3 57.1 55.5 64.7 53.1 48.7 94.2%
SparseVLM(ICML25) 56.4 53.9 60.7 57.3 68.4 55.2 48.1 97.5%
FrameFusion(ICCV25) 56.0 54.8 61.7 57.5 68.2 55.7 48.6 98.1%
GlobalCom2 57.0 55.4 62.6 58.1 69.2 55.8 49.4 99.3%
Ratio=15%, Retain 941 Tokens
FastV(ECCV24) 51.6 48.3 55.0 48.1 51.4 49.4 43.3 85.0%
PDrop(CVPR25) 53.2 47.6 54.7 50.1 58.7 48.7 45.0 87.4%
SparseVLM(ICML25) 52.9 49.7 57.4 53.4 61.0 52.1 47.0 91.2%
GlobalCom2 54.2 52.9 60.7 55.8 66.0 53.3 48.1 95.3%
Method LongVideo. MLVU VideoMME
Overall Short Medium Long
Upper Bound
Qwen2-VL-7B 56.0 60.4 57.6 70.0 54.6 48.2
Ratio=25%
FastV(ECCV24) Out of Memory (OOM)
DyCoke(CVPR25) 50.5 55.2 51.5 62.1 47.1 45.1
GlobalCom2 51.4 55.9 54.3 64.7 50.1 48.2
F.1 Comparisons on LLaVA-NeXT-13B
@@
F.2 Results on LLaVA-OneVision-0.5B
@@
F.3 Results on Video Understanding.
Table 8 presents GlobalCom2’s extension to video understanding with LLaVA-OneVision, demonstrating two key advantages: (i) Superior VideoLLM performance: GlobalCom2 achieves 2.7% average improvement over VideoLLM-specific method DyCoke (Tao et al. 2025) while using fewer vision tokens. (ii) Advantages on long video understanding: At $R=15\%$, GlobalCom2 achieves 3.2 and 3.3 points higher than the second-best method on LongVideoBench and MLVU, demonstrating the effectiveness of our “global-to-local” guided compression design. Table 9 shows GlobalCom2 significantly outperforms DyCoke on Qwen2-VL, confirming its scalability. Additionally, FastV explicitly computes the full attention matrix in the certain layer, causing OOM issues with high token counts and limiting its practical applications.
@@
F.4 Sensitivity Analysis of Hyper-parameters
@@
F.5 Detailed Efficiency Analysis
@@
F.6 More Visualizations of Token Compression
@@
Appendix G Algorithm Illustration